# Learning with Errors over Group Rings Constructed by Semi-direct ${}^{\dagger}$

### 01 Dec 2023 v2

@@ Abstract
Index Terms:
Learning with errors; Group rings; Semi-direct product; Group representations; Lattice-based cryptography; Computational hardness
@@ I Introduction
I-A Latticed-based cryptography
Lattice-based cryptography has been playing a significant role in post-quantum cryptography. It is well-known that traditional cryptographic systems such as RSA and Diffie-Hellman protocol rely on the computational hardness of certain number-theoretic problems. However, Shor [2] introduced a quantum algorithm that can solve integer factorization and discrete logarithm problems in polynomial time, rendering these classical cryptosystems vulnerable to quantum attacks. In contrast, lattice-based cryptosystem has several advantages over such traditional cryptosystems. First, no efficient quantum algorithms have been discovered for lattice problems that are used to construct cryptographic tools for the moment. Consequently, efforts are currently underway to establish post-quantum cryptographic standards, with lattice-based schemes as leading candidates. Second, there is an inherent disparity between the average-case hardness required by cryptography and the worst-case hardness concerned in computational complexity theory. Unlike other types of cryptosystems, lattice-based cryptosystems own security guaranteed by the worst-case hardness of certain lattice problems. Ajtai [3] proposed the first public-key cryptosystem from the short integer solution ($\mathsf{SIS}$) problem whose security is based on the worst-case hardness of well-studied lattice problems. Third, lattice-based cryptography is computationally efficient in comparison with RSA and the Diffie-Hellman protocol, as it only requires linear operations on vectors or matrices modulo relatively small integers, without the need for large number exponentiations.
@@ I-A Latticed-based cryptography
[…]
@@
In [4], Regev initially proved the hardness of the $\mathsf{LWE}$ problem for certain parameters, assuming the hardness of finding short vectors in lattices. Specifically, Regev established a (quantum) reduction from worst-case approximate $\mathsf{SVP}$ (shortest vector problem) and $\mathsf{SIVP}$ (shortest independent vectors problem) for $n$-dimensional lattices to the average-case $\mathsf{LWE}$ problem with the same dimension $n$ as the corresponding lattices. Informally speaking, the reduction suggests that if someone could solve $\mathsf{LWE}$ in polynomial time with non-negligible probability, it would imply the existence of a quantum algorithm capable of solving approximate $\mathsf{SVP}$ or $\mathsf{SIVP}$ in arbitrary lattices with the same dimension as in $\mathsf{LWE}$ problem. Later, Peikert [5] presented a classical reduction from worst-case approximate $\mathsf{SVP}$, but with worse parameters than those in [4].
@@
While Regev’s $\mathsf{LWE}$ enjoys worst-case hardness, it suffers from drawbacks regarding efficiency. It scales poorly with increasing dimension $n$, which makes it impractical for high-dimensional settings. Inspired by the NTRU [6] cryptosystem proposed by Hoffstein et al., which is constructed from algebraic lattices, Stehlé et al. [7] proposed a more efficient variant of the $\mathsf{LWE}$ problem over a cyclotomic ring. Lyubashevsky et al. [1] then presented the reduction from the hard problems in ideal lattices in the worst case to the Ring-$\mathsf{LWE}$ problem. The improved efficiency of this problem stems from the additional algebraic structures of the rings. The utilization of number-theoretic transformation (NTT), a variant of fast Fourier transformation (FFT) also accelerates the multiplication of ring elements in certain cyclotomic rings, contributing to its computational efficiency. Precisely, Ring-$\mathsf{LWE}$ also considers equations of the form $b_{i}=a_{i}\cdot s+e_{i}$ for a secret element $s\in R^{\vee}/qR^{\vee}$ and public uniformly random $a_{i}\in R/qR$, where $q$ is the modulus and $R$ is a commutative ring typically chosen as the ring of algebraic integers of a number field. Despite its additional algebraic structure, no known algorithms for ideal lattices outperform those in general lattices for the moment. Later, Peikert et al. [8] gave a direct quantum reduction from worst-case (ideal) lattice problems to the decision version of (Ring)-$\mathsf{LWE}$. Brakerski et al. [9] introduced Module-$\mathsf{LWE}$, balancing the efficiency and compactness of Ring-$\mathsf{LWE}$ and the hardness of Regev’s plain $\mathsf{LWE}$. Module-$\mathsf{LWE}$ has more flexible parameters than Ring-$\mathsf{LWE}$ and Regev’s standard $\mathsf{LWE}$, making it a reasonable choice for standardization. Many schemes submitted to NIST for post-quantum standardization such as CRYSTALS-Kyber, CRYSTALS-Dilithium, and Saber, rely on the assumed hardness of $\mathsf{LWE}$ variants. Additionally, there are several other algebraically structured $\mathsf{LWE}$ variants, e.g., Polynomial-$\mathsf{LWE}$ [10], Order-$\mathsf{LWE}$ [11], and Middle-Product $\mathsf{LWE}$ [12]. Peikert et al. [13] have developed a unified framework for all proposed $\mathsf{LWE}$ variants (over commutative base rings), which encompasses all prior “algebraic” $\mathsf{LWE}$ variants defined over number fields. This work has led to more simplified and tighter reductions to Ring-$\mathsf{LWE}$.
@@ I-B Our results
Similar to the work in [30], we consider the $\mathsf{LWE}$ problems over group rings. Let $G=\{g_{1},g_{2},\ldots,g_{n}\}$ be a finite group, and let $R$ be a commutative ring, elements in the group ring $R[G]$ are formal sum
@@
where $s_{1}+2s_{2}=n$. Every element $a\in R$ is mapped into $H$ by taking $n$ distinct embeddings of $a$ ($s_{1}$ real embeddings and $2s_{2}$ complex embeddings). It can be easily verified that $H$, with the Hermite inner product, is isomorphic to the Euclidean inner product space $\mathbb{R}^{n}$. Canonical embedding often leads to tighter error rates and enables to obtain more compact systems due to its component-wise multiplication. On the other hand, the coefficient embedding introduced by [7] directly maps the ring elements to real vectors according to its coefficients under a specific basis. In this paper, we primarily use coefficient embedding for the sake of simplicity, following the approach described in [30]. However, we also need to use an extended version of canonical embedding to analyze the generalized $\mathsf{LWE}$ problem due to its connection to irreducible group representations.
@@
In the case of a group ring $\mathbb{Z}[G]$, we can regard it as a free $\mathbb{Z}$-module. All the group elements form a $\mathbb{Z}$-basis of $\mathbb{Z}[G]$ by natural. An element $\sum_{i=1}^{n}r_{i}g_{i}\in\mathbb{R}[G]$ can be represented by an $n$-dimensional vector $(r_{1},r_{2},\ldots,r_{n})\in\mathbb{R}^{n}$. Under this embedding, each ideal of $\mathbb{Z}[G]$ corresponds to a lattice in $\mathbb{R}^{n}$, which is called an ideal lattice. In our paper, we mainly study the relationship between the hard problems in ideal lattices and the $\mathsf{LWE}$ problem over group rings.
@@ I-B1 Avoiding the potential attack
For “hard” problems such as $\mathsf{SVP}$ and $\mathsf{SIVP}$ in certain ideal lattices, several potential attacks have been discovered. Various quantum polynomial-time algorithms have been developed to solve these problems in specific principal ideal lattices as mentioned in [32, 33, 34]. However, it is important to note that the Ring-$\mathsf{LWE}$ problem present in [1] is not under direct threat from the attacks above, as the corresponding lattice problem serves as a lower bound of the security. Regarding the $\mathsf{LWE}$ problem over group rings studied in this paper, it may not be affected by these algorithms for some reasons. On the one hand, the ideals of the group rings may not be principal. On the other hand, these attacks are basically designed for rings with commutative operation. Therefore, the attack can not be directly applied to the scenarios discussed in this paper, where the multiplication operation is not commutative.
@@ I-B2 Why selecting such group rings
Additionally, when performing the multiplication of two elements in the group ring, it is necessary to exploit all irreducible representations of the finite group due to the coefficient embedding. This implies that the highest dimension of irreducible representation cannot be excessively large. It can be verified that the dimension of irreducible representations of $\mathbb{Z}_{m}\ltimes\mathbb{Z}_{n}$ (Type I) is no more than 2, and that the dimension of irreducible representations of $\mathbb{Z}_{n}^{\ast}\ltimes\mathbb{Z}_{n}$ (Type II) is no more than $n-n/p_{1}p_{2}\cdots p_{s}$, where $p_{1},p_{2},\ldots,p_{s}$ are all the distinct prime factors of $s$.
@@ I-B3 Relationship with other $\mathsf{LWE}$ variants
$\mathsf{LWE}$ over group rings involves a non-commutative multiplication operation, unlike the $\mathsf{LWE}$ variants studied in [13]. Consider a group ring $R[G]$, where $G=\{g_{1},g_{2},\ldots,g_{n}\}$ is a finite group of order $n$. In this settings, $g_{1},g_{2},\ldots,g_{n}$ naturally form a $R$-basis of $R[G]$. Let $a,s\in R[G]$, then the product $a\cdot s$ can also be expressed as product of a matrix and a vector with entries from $R$. Precisely, the left multiplication determined by $a$ is an $R$-linear transformation over $\mathbb{R}^{n}$, then $a$ uniquely specifies a unique matrix from $\mathbb{R}^{n\times n}$ under the natural basis, denoted by $\mathcal{M}(a)$. Then the product $a\cdot s$ can be equivalently represented as the product of $\mathcal{M}(a)$ with the coefficient vector of $s$.
$\mathsf{LWE}$ over group rings can also be viewed as a “structured-module” $\mathsf{LWE}$. As stated in [30], the group ring is selected as $\mathbb{Z}[D_{2n}]$, i.e., $\mathbb{Z}[\mathbb{Z}_{2}\ltimes\mathbb{Z}_{n}]$. Let $r,t$ be generators of $\mathbb{Z}_{2}$ and $\mathbb{Z}_{n}$, respectively. Let $s=s_{1}(t)+r\cdot s_{2}(t),a=a_{1}(t)+r\cdot a_{2}(t),b=b_{1}(t)+r\cdot b_{2}(t)\in\mathbb{Z}[D_{2n}]$, where $s_{1}(t),s_{2}(t),a_{1}(t),a_{2}(t),b_{1}(t),b_{2}(t)\in\mathbb{Z}[t]/\langle t^{n}-1\rangle,$ then the equation $b\approx a\cdot s$ can be expressed in the matrix form as follows:
$\left(\begin{array}[]{c}b_{1}\\ b_{2}\end{array}\right)\approx M\cdot\left(\begin{array}[]{c}s_{1}\\ s_{2}\end{array}\right).$
Here, $M$ is a structured $2\times 2$ matrix with elements in $\mathbb{Z}[t]/\langle t^{n}-1\rangle$ all of which can be computed from $s_{1}(t),s_{2}(t),a_{1}(t),a_{2}(t)$ efficiently. It is worth noting that the matrix $M$ is structured, since only two elements ($a_{1}(t)$ and $a_{2}(t)$) of the four need to be sampled randomly, while in the situation of unstructured-module $\mathsf{LWE}$, all entries in $M$ are sampled randomly as in [37]. That’s the reason why the $\mathsf{LWE}$ over group rings can be viewed as a “structured-module” $\mathsf{LWE}$.
@@ I-B4 Proof of reductions
The first reduction is from a worst-case lattice problem to the search $\mathsf{GR}\mbox{-}\mathsf{LWE}$ problem. The reduction is based on the work of [1] and [13]. It can be applied to any group ring that satisfies certain mild properties. Precisely, the reduction applies to those group rings satisfying if the ideal of the group ring has an inverse ideal, then the inverse ideal and the dual ideal are essentially equivalent, up to a certain permutation of the coefficients. Since not any ideal of the group rings is invertible, this reduction only considers those invertible ones. With the help of the search $\mathsf{GR}\mbox{-}\mathsf{LWE}$ oracle, we are capable of sampling from discrete Gaussian distributions with narrower and narrower widths in (invertible) ideal lattices through iterative steps. This process continues until the Gaussian parameter reaches our expectations. Consequently, we can obtain “short” vectors in the lattice except with negligible probability.
The second one is a direct reduction from a (worst-case) lattice problem to the (average-case) decision $\mathsf{GR}\mbox{-}\mathsf{LWE}$ problem. The procedure closely follows the approaches outlined in [8] and [30]. The coefficient embedding and generalized canonical embedding both play crucial roles in this reduction. We establish the hardness by providing the reduction from solving bounded distance decoding ($\mathsf{BDD}$), the problem of finding the closest lattice vector to a target $\mathbf{t}\in\mathbb{R}^{n}$ that is sufficiently close to a lattice $\mathcal{L}$, to the decision $\mathsf{GR}$$\mathsf{LWE}$ problem. During the reduction process, the decision $\mathsf{LWE}$ oracle behaves like an oracle with a “hidden center” defined by [8]. To see how the oracle works for reduction, [4] states that one can transform an $\mathsf{BDD}$ instance into an $\mathsf{LWE}$ sample whose secret corresponds to the closest lattice vector $\mathbf{v}\in\mathcal{L}$ to $\mathbf{t}$. With a suitable oracle for decision $\mathsf{LWE}$ and by incrementally moving $\mathbf{t}$ towards the closest lattice vector $\mathbf{v}$ by carefully measuring the behavior of the oracle, we can detect the distance between the moving point $\mathbf{t}$ and $\mathcal{L}$ by monitoring the acceptance probability of the oracle, which is a monotonically decreasing function of $\operatorname{dist}(\mathbf{t},\mathcal{L})$. This makes it possible to solve $\mathsf{BDD}$ by repeatedly perturbing $\mathbf{t}$ to a new point $\mathbf{t}^{\prime}$, and testing whether the new point is significantly closer to the lattice. For further details, we refer to Section IV.

### 10 Mar 2026 v3

# Learning with Errors over Group Rings Constructed by Semi-direct Product

@@ I Introduction
I-A Lattice-based cryptography
Lattice-based cryptography has been playing a significant role in post-quantum cryptography. It is well-known that traditional cryptographic systems such as RSA and Diffie-Hellman protocol rely on the computational hardness of certain number-theoretic problems. However, Shor [2] introduced a quantum algorithm that can solve integer factorization and discrete logarithm problems in polynomial time, rendering these classical cryptosystems vulnerable to quantum attacks. In contrast, lattice-based cryptosystems have several advantages over such traditional cryptosystems. First, no efficient quantum algorithms have been discovered for lattice problems that are used to construct cryptographic tools for the moment. Consequently, efforts are currently underway to establish post-quantum cryptographic standards, with lattice-based schemes as leading candidates. Second, there is an inherent disparity between the average-case hardness required by cryptography and the worst-case hardness concerned in computational complexity theory. Unlike other types of cryptosystems, lattice-based cryptosystems own security guaranteed by the worst-case hardness of certain lattice problems when their parameters are suitably chosen. Ajtai [3] proposed the first public-key cryptosystem from the short integer solution ($\mathsf{SIS}$) problem whose security is based on the worst-case hardness of well-studied lattice problems. Third, lattice-based cryptography is computationally efficient in comparison with RSA and the Diffie-Hellman protocol, as it only requires linear operations on vectors or matrices modulo relatively small integers, without the need for large number exponentiations.
@@ I-A Lattice-based cryptography
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product in $\mathbb{R}^{n}$. The $\mathsf{LWE}$ problem asks to recover the secret vector $\mathbf{s}$ given $\mathsf{LWE}$ samples (search version) or distinguish $\mathsf{LWE}$ samples from uniformly random samples in $\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$ with a non-negligible advantage within polynomial time (decision version).
@@
In [4], Regev initially proved the hardness of the $\mathsf{LWE}$ problem for certain parameters, assuming the hardness of finding short vectors in lattices. Specifically, Regev established a (quantum) reduction from worst-case (decisional) approximate $\mathsf{SVP}$ (shortest vector problem) and $\mathsf{SIVP}$ (shortest independent vectors problem) for $n$-dimensional lattices to the average-case $\mathsf{LWE}$ problem with the same dimension $n$ as the corresponding lattices. Informally speaking, the reduction suggests that if someone could solve $\mathsf{LWE}$ in polynomial time with non-negligible probability, it would imply the existence of a quantum algorithm capable of solving approximate $\mathsf{SVP}$ or $\mathsf{SIVP}$ in arbitrary lattices with the same dimension as in $\mathsf{LWE}$ problem. Later, Peikert [5] presented a classical reduction from worst-case approximate $\mathsf{SVP}$, but with worse parameters than those in [4].
@@
While Regev’s $\mathsf{LWE}$ enjoys worst-case hardness, it suffers from drawbacks regarding efficiency. It scales poorly with increasing dimension $n$, which makes it impractical for high-dimensional settings. Inspired by the NTRU [6] cryptosystem proposed by Hoffstein et al., which is constructed from algebraic lattices, Stehlé et al. [7] proposed a more efficient variant of the $\mathsf{LWE}$ problem over a cyclotomic ring. Lyubashevsky et al. [1] then presented a (quantum) worst-case-to-average-case reduction from ideal lattice problems in the ring of integers of certain number fields (notably cyclotomic/Galois ones) to the Ring-$\mathsf{LWE}$ problem. The improved efficiency of this problem stems from the additional algebraic structures of the rings. The utilization of number-theoretic transformation (NTT), a variant of fast Fourier transformation (FFT) also accelerates the multiplication of ring elements in certain cyclotomic rings, contributing to its computational efficiency. Precisely, Ring-$\mathsf{LWE}$ also considers equations of the form $b_{i}=a_{i}\cdot s+e_{i}$ for a secret element $s\in R^{\vee}/qR^{\vee}$ and public uniformly random $a_{i}\in R/qR$, where $q$ is the modulus, $R$ is a commutative ring typically chosen as the ring of algebraic integers of a number field and $R^{\vee}$ is the dual ideal of $R$. Although some algorithms are more efficient on ideal lattices than on general lattices due to their additional algebraic structures, none are currently known to threaten well-designed cryptographic constructions. In particular, no algorithms are known that significantly outperform the best general lattice algorithms in attacking such schemes. Later, Peikert et al. [8] gave a direct quantum reduction from worst-case (ideal) lattice problems to the decision version of (Ring)-$\mathsf{LWE}$. Brakerski et al. [9] introduced Module-$\mathsf{LWE}$, balancing the efficiency and compactness of Ring-$\mathsf{LWE}$ and the hardness of Regev’s plain $\mathsf{LWE}$. Module-$\mathsf{LWE}$ has more flexible parameters than Ring-$\mathsf{LWE}$ and Regev’s standard $\mathsf{LWE}$, making it a reasonable choice for standardization. Many schemes submitted to NIST for post-quantum standardization such as CRYSTALS-Kyber, CRYSTALS-Dilithium, and Saber, rely on the assumed hardness of $\mathsf{LWE}$ variants. Additionally, there are several other algebraically structured $\mathsf{LWE}$ variants, e.g., Polynomial-$\mathsf{LWE}$ [10], Order-$\mathsf{LWE}$ [11], and Middle-Product $\mathsf{LWE}$ [12]. Peikert et al. [13] have developed a unified framework for all proposed $\mathsf{LWE}$ variants (over commutative base rings), which encompasses all prior “algebraic” $\mathsf{LWE}$ variants defined over number fields. This work has led to more simplified and tighter reductions to Ring-$\mathsf{LWE}$.
@@ I-B Our results
Similar to the work in [30], we consider the $\mathsf{LWE}$ problems over group rings. Let $G=\{g_{1},g_{2},\ldots,g_{n}\}$ be a finite group, and let $R$ be a commutative ring, elements in the group ring $R[G]$ are formal sums
@@
where $s_{1}+2s_{2}=n$. Every element $a\in R$ is mapped into $H$ by taking $n$ distinct embeddings of $a$ ($s_{1}$ real embeddings and $2s_{2}$ complex embeddings). It can be easily verified that $H$, with the Hermite inner product, is isomorphic to the Euclidean inner product space $\mathbb{R}^{n}$. Canonical embedding often leads to tighter error rates and enables to obtain more compact systems due to its component-wise multiplication. On the other hand, the coefficient embedding used in [7] directly maps the ring elements to real vectors according to its coefficients under a specific basis. In this paper, we primarily use coefficient embedding for the sake of simplicity, following the approach described in [30]. However, we also need to use an extended version of canonical embedding to analyze the generalized $\mathsf{LWE}$ problem due to its connection to irreducible group representations.
@@
In the case of a group ring $\mathbb{Z}[G]$, we can regard it as a free $\mathbb{Z}$-module. All the group elements form a $\mathbb{Z}$-basis of $\mathbb{Z}[G]$ inherently. An element $\sum_{i=1}^{n}r_{i}g_{i}\in\mathbb{R}[G]$ can be represented by an $n$-dimensional vector $(r_{1},r_{2},\ldots,r_{n})\in\mathbb{R}^{n}$. Under this embedding, each ideal of $\mathbb{Z}[G]$ corresponds to a lattice in $\mathbb{R}^{n}$, which is called an ideal lattice. In our paper, we mainly study the relationship between the hard problems in ideal lattices and the $\mathsf{LWE}$ problem over group rings.
@@ I-B1 Avoiding the potential attack
For “hard” problems such as $\mathsf{SVP}$ and $\mathsf{SIVP}$ in certain ideal lattices, several potential attacks have been discovered. Various quantum polynomial-time algorithms have been developed to solve these problems in specific principal ideal lattices as mentioned in [32, 33, 34]. Some studies generalized the SVP algorithms to non-principal ideal lattices such as mentioned in [35]. However, it is important to note that the Ring-$\mathsf{LWE}$ problem present in [1] is not under direct threat from the attacks above, as the corresponding lattice problem serves as a lower bound of the security. Regarding the $\mathsf{LWE}$ problem over group rings studied in this paper, it may not be affected by these algorithms for some reasons. On the one hand, the ideals of the group rings may not be principal. On the other hand, these attacks are basically designed for rings with commutative operation. Therefore, the attack can not be directly applied to the scenarios discussed in this paper, where the multiplication operation is not commutative.
@@ I-B2 Why selecting such group rings
Additionally, when performing the multiplication of two elements in the group ring, it is necessary to exploit all irreducible representations of the finite group due to the coefficient embedding. This implies that the highest dimension of irreducible representation cannot be excessively large. It can be verified that the dimension of irreducible representations of $\mathbb{Z}_{m}\ltimes\mathbb{Z}_{n}$ (Type I) is no more than 2, and that the dimension of irreducible representations of $\mathbb{Z}_{n}^{\ast}\ltimes\mathbb{Z}_{n}$ (Type II) is no more than $n-n/p_{1}p_{2}\cdots p_{s}$, where $p_{1},p_{2},\ldots,p_{s}$ are all the distinct prime factors of $n$.
@@ I-B3 Relationship with other $\mathsf{LWE}$ variants
$\mathsf{LWE}$ over group rings involves a non-commutative multiplication operation, unlike the $\mathsf{LWE}$ variants studied in [13], where commutative algebraic systems are mainly concerned. Consider a group ring $R[G]$, where $G=\{g_{1},g_{2},\ldots,g_{n}\}$ is a finite group of order $n$. In this settings, $g_{1},g_{2},\ldots,g_{n}$ naturally form a $R$-basis of $R[G]$. Let $a,s\in R[G]$, then the product $a\cdot s$ can also be expressed as product of a matrix and a vector with entries from $R$. Precisely, the left multiplication determined by $a$ is an $R$-linear transformation over $\mathbb{R}^{n}$, then $a$ uniquely specifies a unique matrix from $\mathbb{R}^{n\times n}$ under the natural basis, denoted by $\mathcal{M}(a)$. Then the product $a\cdot s$ can be equivalently represented as the product of $\mathcal{M}(a)$ with the coefficient vector of $s$.
$\mathsf{LWE}$ over group rings can also be viewed as a “structured-module” $\mathsf{LWE}$. As stated in [30], the group ring is selected as $\mathbb{Z}[D_{2n}]$, i.e., $\mathbb{Z}[\mathbb{Z}_{2}\ltimes\mathbb{Z}_{n}]$. Let $r,t$ be generators of $\mathbb{Z}_{2}$ and $\mathbb{Z}_{n}$, respectively. Let $s=s_{1}(t)+s_{2}(t)\cdot r,a=a_{1}(t)+a_{2}(t)\cdot r,b=b_{1}(t)+b_{2}(t)\cdot r\in\mathbb{Z}[D_{2n}]$, where $s_{1}(t),s_{2}(t),a_{1}(t),a_{2}(t),b_{1}(t),b_{2}(t)\in\mathbb{Z}[t]/\langle t^{n}-1\rangle,$ then the equation $b\approx s\cdot a$ can be expressed in the matrix form as follows:
$\left(\begin{array}[]{c}b_{1}\\ b_{2}\end{array}\right)\approx\mathbf{M}\cdot\left(\begin{array}[]{c}s_{1}\\ s_{2}\end{array}\right).$
Here, $\mathbf{M}$ is a structured $2\times 2$ matrix with elements in $\mathbb{Z}[t]/\langle t^{n}-1\rangle$ all of which can be computed from $a_{1}(t),a_{2}(t)$ efficiently. It is worth noting that the matrix $\mathbf{M}$ is structured, since only two elements ($a_{1}(t)$ and $a_{2}(t)$) of the four need to be sampled randomly, while in the situation of unstructured-module $\mathsf{LWE}$, all entries in $\mathbf{M}$ are sampled randomly as in [38]. That’s the reason why the $\mathsf{LWE}$ over group rings can be viewed as a “structured-module” $\mathsf{LWE}$. Compared with the classical module $\mathsf{LWE}$, to produce the same number of pseudorandom $\mathsf{LWE}$ samples, the additional structure can reduce the cost of uniform randomness for sampling $\mathbf{M}$ by a factor of $1/2$.
@@ I-B4 Proof of reductions
The first reduction is from a worst-case lattice problem to the search $\mathsf{GR}\mbox{-}\mathsf{LWE}$ problem. The reduction is based on the work of [1] and [13]. It can be applied to any group ring that satisfies certain mild properties. Precisely, the reduction applies to those group rings satisfying if the ideal of the group ring has an inverse ideal, then the inverse ideal and the dual ideal are essentially equivalent, up to a certain permutation of the coefficients. Since not every ideal of the group rings is invertible, this reduction only considers those invertible ones. With the help of the search $\mathsf{GR}\mbox{-}\mathsf{LWE}$ oracle, we are capable of sampling from discrete Gaussian distributions with narrower and narrower widths in (invertible) ideal lattices through iterative steps. This process continues until the Gaussian parameter reaches our expectations. Consequently, we can obtain “short” vectors in the lattice except with negligible probability.
The second one is a direct reduction from a (worst-case) lattice problem to the (average-case) decision $\mathsf{GR}\mbox{-}\mathsf{LWE}$ problem. The procedure closely follows the approaches outlined in [8] and [30]. The coefficient embedding and generalized canonical embedding both play crucial roles in this reduction. We establish the hardness by providing the reduction from solving bounded distance decoding ($\mathsf{BDD}$), the problem of finding the closest lattice vector to a target $\mathbf{t}\in\mathbb{R}^{n}$ that is sufficiently close to a lattice $\mathcal{L}$, to the decision $\mathsf{GR}\mbox{-}\mathsf{LWE}$ problem. During the reduction process, the decision $\mathsf{LWE}$ oracle behaves like an oracle with a “hidden center” defined by [8]. To see how the oracle works for reduction, [4] states that one can transform an $\mathsf{BDD}$ instance into an $\mathsf{LWE}$ sample whose secret corresponds to the closest lattice vector $\mathbf{v}\in\mathcal{L}$ to $\mathbf{t}$. With a suitable oracle for decision $\mathsf{LWE}$ and by incrementally moving $\mathbf{t}$ towards the closest lattice vector $\mathbf{v}$ by carefully measuring the behavior of the oracle, we can detect the distance between the moving point $\mathbf{t}$ and $\mathcal{L}$ by monitoring the acceptance probability of the oracle, which is a monotonically decreasing function of $\operatorname{dist}(\mathbf{t},\mathcal{L})$. This makes it possible to solve $\mathsf{BDD}$ by repeatedly perturbing $\mathbf{t}$ to a new point $\mathbf{t}^{\prime}$, and testing whether the new point is significantly closer to the lattice. For further details, we refer to Section IV.